Reg 1A Reg 1B Req ЗА Req 3B Compute the break-even point in dollar sales for April. (Round intermediate percentage calculatio| Regia Req 1A Reg 1B Reg 18 Req3a Reg 3A | Regae | Reg 3B Would the break-even point in May be higher or lower than the breaCarbex, Inc., produces cutlery sets out of high-quality wood and steel. The company makes a Standard set and a Deluxe set andComplete this question by entering your answers in the tabs below. Reg 1A Reg 1B Req ЗА Reg 3B Prepare contribution format inReg 1A Reg 1B Reg 3A Reg 3B Prepare contribution format income statements for May. (Round your Percentage answers to 1 deci
Answer #1

1-a April

Standard Deluxe Total
Amount % Amount % Amount %
Sales Revenue $    3,47,800 100.00% $          2,40,300 100.00% $ 5,88,100 100.00%
Variable Expenses
Variable Production costs $    1,03,400 29.73% $             99,900 41.57% $ 2,03,300 34.57%
Sales commission $       76,516 22.00% $             52,866 22.00% $ 1,29,382 22.00%
Total Variable Expenses $    1,79,916 51.73% $          1,52,766 63.57% $ 3,32,682 56.57%
Contribution Margin $    1,67,884 48.27% $             87,534 36.43% $ 2,55,418 43.43%
Fixed Expenses
Advertising $ 1,12,000
Depreciation $    23,800
Administrative $    66,500
Total Fixed Expenses $ 2,02,300
Net Operating Income $    53,118

1-b May

1700 5700
Standard Deluxe Total
Amount % Amount % Amount %
Sales Revenue $    1,25,800 100.00% $          5,07,300 100.00% $ 6,33,100 100.00%
Variable Expenses
Variable Production costs $       37,400 29.73% $          2,10,900 41.57% $ 2,48,300 39.22%
Sales commission $       27,676 22.00% $          1,11,606 22.00% $ 1,39,282 22.00%
Total Variable Expenses $       65,076 51.73% $          3,22,506 63.57% $ 3,87,582 61.22%
Contribution Margin $       60,724 48.27% $          1,84,794 36.43% $ 2,45,518 38.78%
Fixed Expenses
Advertising $ 1,12,000
Depreciation $    23,800
Administrative $    66,500
Total Fixed Expenses $ 2,02,300
Net Operating Income $    43,218

2-a Break Even point (April) = Fixed expenses / Weighted Contribution margin per unit
= $202300 / $34.52 = 5862 units
Weighted average contribution margin = $255418 / 7400 = $34.52
Standard = 5862 x 4700/7400 = 3723 units x $74 = $275502
Deluxe = 5862 x 2700/7400 = 2139 units x $89 = $190282
Total Break even point = $465784

2-b  Break Even point (April) = Fixed expenses / Weighted Contribution margin per unit
= $202300 / $33.18 = 6098 units
Weighted average contribution margin = $245518 / 7400 = $33.18
Standard = 6098 x 1700/7400 = 1401 units x $74 = $103674
Deluxe = 6098 x 5700/7400 = 4697 units x $89 = $418033
Total Break even point = $521707

Break even point will be higher since sales of standard units with higher contribution is lower